A total of 285 tickets were sold for the school play. The number of student tickets sold was two times the number of adult tickets sold . How many adult tickets were sold?

Answer by Earlsdon(6294)

Let S = the number of student tickets sold and A = the number of adult tickets sold.
From the problem, you can write:
1) S+A = 285 "A total of 285 (student plus adult) tickets were sold..."
2) S = 2A "The number of student tickets (S) sold was two times the number of adult tickets (A) sold."
Substitute equation 2) (S = 2A) for S in equation 1).
(2A+A) = 285 Simplify.
3A = 285 Divide both sides by 3.
A = 95 This is the number of adult tickets sold.
From equation 2)...
S = 2A
S = 2(95)
S = 190 This is the number of student tickets sold.